RC meeting and Run Plan, Thursday Sept. 28, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for major target positions are on the Control Room whiteboard.)

Status Update

Overall Summary

No beam all of yesterday (09-27) DAY or SWING shifts. OWL received enough beam to check the harp scans, but it dropped as data was being taken on the carbon hole. The issue that stopped the beam then is still ongoing, and will likely not be fixed before the end of DAY shift today.

Overnight, there were again issues with the newer laser computer going into hibernation and turning off the associated lasers. A temporary workaround for this has been found, but any incoming target operators need to ensure that the computer stays 'awake'.

The DAY shift today has been taking cosmics data for BBCAL calibrations.

A Lenovo power supply has been brought out of the hall (and has been cleared by RadCon). If this is yours, it's in the Counting House.

Accelerator status

  • Yet more issues with the PSS in the BSY. MCC are trying to get this fixed; or at least, to find a workaround. This is expected to be done some time around the start of SWING, but this is subject to change.
  • No maintenance day on the MCC white board (until 10/10/2023).

Data collected :

  • Accumulated corrected charge collected: 0C (no beam = no charge...)

Hall A Subsystems Status


  • Latest target polarization: 42.3% [11:03 am]
  • There are still brief (~1 minute or less) spikes in the pumping cell temperature occurring not infrequently (see https://logbooks.jlab.org/entry/4188853). These seem to return back to normal temperatures very quickly, and only trigger the pumping cell alarms.


  • BBHodo: reverted changes to HV following tests over the weekend.
  • HCAL: Jiwan made an escorted access to replace a fADC module, but this was unsuccessful.

Hall access requests

  • HCAL (Jiwan): Attempt to repeat fADC replacement if the module is ready.
  • If anyone else wants to make an access, let me know! There is no planned maintenance day on Tuesday 3rd, so any accesses need to be opportunistic.

Shift schedule

  • Still need to fill open slots
  • Open slots on Saturday 30th: DAY (TO) and SWING (SL).


Please sign up for shifts if there are any open!

Short Term Run plan

  • Beam recovery to the hall
  • Production run on polarized He3 target @35uA beam current.
    • If the beam and target are stable at 35uA, call the RC. We can then go up to 40uA.

==> List of items in case He3 production running isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)